tically you can open Assignment 1 here 4. A student calibrates a 25-mL pipette by measuring the volume of the water it can transfer. The reading data are shown as follows (unit in mL): 24.75, 24.83, 24.87, 24.91, and 25.32 (a) Is there an outlier at the 95% confidence level? (b) Calculate the coefficient of variation based on the results from part (a) (c) Suppose the standard deviation of the liquid volume is 0.04 mL, how many replicate measurements would the student need to get 95% confidence interval limited to 0.02 mL? (d) The pipette is well calibrated if the calibration results show no significant difference with the calibration results done by the manufactory at the 95% confidence level. The standard results are shown here Standard results No. of sample Mean Standard deviation 24.97 ml 0.071 ml Based on the results from part (a), is the pipette well calibrated? 5. The blood glucose level of a person can be measured by spectrometry for clinical use. The absorbance intensity of 6 standard solutions with known glucose concentration is shown below Conc. of glucose (M) Absorbance 0 2x10 0.189 4x10 0.412 6x10 0.672 8x10 0.818 lx10 1.104 (a) Calculate the equation of the best-fit straight line for the standard glucose solution using the
Answer #1

data

mean = 24.936

sd = 0.2227

95% confidence interval

(24.936 - 2*0.2227 , 24.936 + 2*0.2227)

=( 24.4906 , 25.3814 )

since each value is present in confidence interval

there is not outlier

b)

coefficient of variation = sd/mean =

0.2227 / 24.936

= 0.00893

c)

e = 0.02

n = (z * sd/e)^2

=(1.96 * 0.2227/0.02)^2

= 476.313

=477

d)

95% confidence interval of manufactory = (24.97 - 2 * 0.071 , 24.97 + 2*0.071)

= ( 24.828 , 25.112 )

since 24.75 and 25.32 lies outside this CI

pipette is not well calibrated